# Strong electron-electron interactions in a dilute weakly-localized metal near a metal-to-insulator transition

### v1

@@ Abstract
Si MOSFET MC_silicon_inversion_layers 48 252 0.4-0.8$\times$10-4 10 10-20
QW 2DHG Shashkin_2005 26 61 0.4-3.8$\times$10-4 0.5 10-20
@@
[…]
@@
Figure 2: Typical magneto-conductance data. (a) At temperature $T=115{\rm}$ mK, the magnetic field is set parallel to the plane of the $\delta$-layer (blue points - with no dependence on the relative orientation to the current) and perpendicular to it (green points). The black points are the difference between in- and out-of-plane magneto-conductance per square, the solid black line being a fit with weak-localization corrections, Eq. (1) minus Eq. (2). The orange data (right y-axis) is a Hall measurement ($R_{xy}$) taken at 1.75 K, fitted linearly by the solid line. (b) Difference between out-of-plane and in-plane magneto-conductance per square for various temperatures. The solid lines are fits with Eq. (1) minus Eq. (2). (c) Conductance per square versus temperature in black, displaying a logarithmic dependence as expected for weak localization down to $T_{\text{c}}$approx 300 mK. The coherence length $\ell_{\phi}$ obtained from the fits in (b) is shown in orange. It follows a power law $T^{-0.31}$ down to $T_{\text{c}}$. All data in (a), (b), and (c) are obtained from a sample of electron density $n=(1.18\pm 0.01)\times 10^{13}$ cm-2.
@@
The weak localization effect calculated in Eq. (1) stems from an expansion in $(k_{\text{F}}\ell)^{-1}$, where $k_{\text{F}}$ is the Fermi momentum of the electrons, which is proportional to the square root of the dopant density. For samples with different densities, the weak localization correction depends primarily on $k_{\text{F}}\ell$, which is consistent with the monotonic dependence of the perpendicular magneto-conductance as a function of $k_{\text{F}}\ell$, as shown in Fig. 3(a) and Fig. 3(b) for $B=2~T$ (see the supplemental section I for information on the variability of $\ell$ and $k_{\text{F}}$ across samples). The electron-electron interaction effects depend on the ratio of Coulomb interaction and kinetic energy, which is proportional to the ionization energy divided by the hopping integral (the donors’ wavefunction overlap) and thus is stronger for lower carrier densities. For in-plane magnetic fields, when the $\delta$-layer is thin and/or dilute enough, the electron-electron interaction contributions dominate over the weak localization term. This is the case for our arsenic $\delta$-layers, as is visible in Fig. 3(a). This regime of thin and dilute half-filled layers has not been achieved in $\delta$-layers before.

### v2

@@ Abstract
[…]
@@
Figure 1: Experimental setup. (a) Hall-bar geometry of the contacts allowing for 4-point resistance measurements. (b) Schematic of the sample composition. The donor layer is encapsulated in the silicon lattice at a depth $d\sim$ 20 nm below the surface, covering a width $\delta$ between 0.4 nm and 1.8 nm having a 2D electron density $n_{\text{2D}}=(1$$19)\times 10^{13}$cm-2.
@@
Figure 2: Typical magneto-conductance data. (a) At temperature $T=115{\rm}$ mK, the magnetic field is set parallel to the plane of the $\delta$-layer (blue points - with no dependence on the relative orientation to the current) and perpendicular to it (green points). The black points are the difference between in- and out-of-plane magneto-conductance per square, the solid black line being a fit with weak-localization corrections, Eq. (1) minus Eq. (2). The orange data (right y-axis) is a Hall measurement ($R_{xy}$) taken at 1.75 K, fitted linearly by the solid line. (b) Difference between out-of-plane and in-plane magneto-conductance per square for various temperatures. The solid lines are fits with Eq. (1) minus Eq. (2). (c) Conductance per square versus temperature in black, displaying a logarithmic dependence as expected for weak localization down to $T_{\text{c}}$ $\approx$ 300 mK. The coherence length $\ell_{\phi}$ obtained from the fits in (b) is shown in orange. It follows a power law $T^{-0.31}$ down to $T_{\text{c}}$. All data in (a), (b), and (c) are obtained from a sample of electron density $n=(1.18\pm 0.01)\times 10^{13}$ cm-2.
@@
The weak localization effect calculated in Eq. (1) stems from an expansion in $(k_{\text{F}}\ell)^{-1}$, where $k_{\text{F}}$ is the Fermi momentum of the electrons, which is proportional to the square root of the dopant density. For samples with different densities, the weak localization correction depends primarily on $k_{\text{F}}\ell$, which is consistent with the monotonic dependence of the perpendicular magneto-conductance as a function of $k_{\text{F}}\ell$, as shown in Fig. 3(a) and Fig. 3(b) for B=2TB=2~T (see the supplemental section I for information on the variability of $\ell$ and $k_{\text{F}}$ across samples). The electron-electron interaction effects depend on the ratio of Coulomb interaction and kinetic energy, which is proportional to the ionization energy divided by the hopping integral (the donors’ wavefunction overlap) and thus is stronger for lower carrier densities. For in-plane magnetic fields, when the $\delta$-layer is thin and/or dilute enough, the electron-electron interaction contributions dominate over the weak localization term. This is the case for our arsenic $\delta$-layers, as is visible in Fig. 3(a). This regime of thin and dilute half-filled layers has not been achieved in $\delta$-layers before.